A drone has a frequency of 2.71 GHz (gigahertz). What wavelength is associated with this frequency? 


1
Expert's answer
2021-09-26T12:06:19-0400

The wavelength is


"\lambda=\frac cf=\frac{3·10^8}{2.71·10^9}=0.111\text{ m}."